OpenStudy (anonymous):

Alicia can row 7 miles downstream in the same time it takes her to row 4 miles upstream. She rows downstream 5 miles/hour faster than she rows upstream. Find Alicia’s rowing rate each way

OpenStudy (anonymous):

Sove the following for r: \[\frac{7}{r+5}= \frac{4}{r}\] She rows downstream at 11 and 2/3 mph and upstream at 6 and 2/3 mph.